OpenStudy (anonymous):

Will give medals (: Using synthetic division, what is the quotient: (x3 + 7x2 + 5x - 20) ÷ (x + 3) ? (3 points) Answer x2 - 4x + 7 x2 + 4x + 7 x2 - 4x - 7 + r(1) x2 + 4x - 7 + r(1)

OpenStudy (campbell_st):

 so \[x^3 + 7x^2 + 5x -20 = (x+3)(x^2 + 4x -7) r = 1\]